BIMOLECULAR REACTIONS OF A DIFFERENT COLOR: CH_{3}D + CHLORINE WITH VARIED PHO-TOLYSIS WAVELENGTHS

\underline{ANDREW E. BERKE}, CHRISTOPHER J. ANNESLEY, and F. FLEMING CRIM, Chemistry Department, University of Wisconsin - Madison, Madison, Wisconsin 53706.

We examine the effect varying the collision energy has on the bimolecular reaction of CH_{3}D and chlorine. A Raman shifting cell pumped at 355 nm and filled with either hygrogen or methane gas provides, in discrete steps, light to photolyze Cl_{2}. This imparts bewteen 600 and 2000 cm^{−1} of collision energy to our system. By also adding C-H vibrational overtone excitation, around 6000 cm^{−1}, we can compare to previous, fixed-photolysis energy studies from our reseach group. We seek to elucidate the role translational energy (collision energy) plays in this reactive system.